Project

General

Profile

NDOS Commissioning Meeting 2011/06/06

Attendance: Jon, Art, Mark, Gary, Andrew, Chad, Zeb, Alena, Jarek, Sue, Mark, Kurt, Ron, Gavin, Leon, Athans

Agenda:

  1. NDOS Status and Channel Recovery Plans
  2. DAQ/Electronics Update & Plans
  3. DCS Update & Plans
  4. AOB (All)

NDOS Status and Channel Recovery Plans

  • Jon:
    • uptime (run>30 minutes long and all 9 dcms in the readout) during the last week was ~80%
    • active channel fraction this past week:
      • whole-detector hovered around 70%
      • fiducial volume hovered around 75%
  • Art: the DAQ ran for 2.5 days with no glitches over the weekend
  • Mark: Steve Dixon suggested last week that we move the HVAC system in the NDSB to a closed-system instead of pulling air in from the outside.

DAQ/Electronics Update & Plans

  • Andrew: 2nd timing unit in MINOS is hooked up and reporting beam spills. Last week a test was done to look at difference between MINOS building timing and Feynman building timing. Expected to see a small time difference, but instead found a 442.19 us time difference (late). This seems non-physical, and so the test will need to be done again. This will happen today.

DCS Update & Plans

  • Ron:
    • IOCs have been running for ~1.5 weeks with issue that, when a DSO scan is done, there is a conflict. So for now, at the beginning of a DSO scan, the IOCs are killed. IOCs are started when DCMs are rebooted (eg, after a DSO scan is done). Priority for this week is to archive temperatures to dB.
    • We have a new person from UVa (Martin) who will work on the display/interface.
    • When can we run a test?
  • Jon:
    • have to get archiving worked out
    • need an interface to enable/disable cooling
    • need tables of set points (Leon), so we only cool the ones that we know operate ok when cooled. Need to get list from Mat and Denis.
  • Mark: what's the procedure for dealing with a fault?
  • Athans: we have not yet decided.
  • Andrew: the hardware protects itself.
  • Goal is to run full test on NDOS by Friday morning.

AOB (All)